# Second- and third-order properties of multidimensional Langevin equations

### v14

@@ 8.4. Dynamics on long time-scales
(391) $\qquad\quad{}+\left\langle\frac{\mathrm{d}[\widetilde{x}^{i},\widetilde{x}^{j}](\tau)}{\mathrm{d}\tau}\mid\widetilde{\mathbf{x}}(0),\widetilde{\mathbf{v}}(0)\right\rangle.$
We now investigate the contributions of the various terms in Eq. (391) to $\langle(\widetilde{x}^{i}(\tau)-\widetilde{x}^{i}(0))(\widetilde{x}^{j}(\tau)-\widetilde{x}^{j}(0))\mid\widetilde{\mathbf{x}}(0),\widetilde{\mathbf{v}}(0)\rangle$ for $\tau\ll\mathcal{O}(\mu)$ [^22] If we immediately put $\tau\sim\mathcal{O}(\mu)$, then by Eqs. (171) and (172), we will introduce a cubic term, which significantly complicates matters.. We do this by multiplying by 1, $\widetilde{\mathbf{x}}(0)$, or $\widetilde{\mathbf{v}}(0)$, taking the expectation and integrating. First, we address the multiplication by 1. The last line of Eq. (391) contributes $\mathcal{O}(\mu^{-1}\tau)$, while the first three lines of Eq. (391) contribute $\mathcal{O}((\mu^{-1}\tau)^{2})$. The remaining terms contribute $\mathcal{O}(\mu^{0})a\mathcal{O}(a,b)\tau$. Thus, only the last line of Eq. (391) survives. Now, we address the multiplication by $\widetilde{x}^{m}(0)$. The last line of Eq. (391) contributes $\mathcal{O}(\mu^{-1}\tau)b$ (to be calculated later), while the first three lines of Eq. (391) contribute $\mathcal{O}((\mu^{-1}\tau)^{2})\mathcal{O}(a,b)$. The fourth line of Eq. (391) contributes:
(392) $-2(\widehat{\mathbf{A}}_{\mathbf{v}}^{-1})^{i}_{i^{\prime}}[(a_{xv})^{i^{\prime}}_{kl}+\mathcal{O}(\mu^{-1})a]\left\{{\langle\widetilde{x}^{k}\widetilde{x}^{m}\rangle[\langle\widetilde{x}^{j}\widetilde{v}^{l}\rangle\tau+\mathcal{O}(\mu^{-2})]+\mathcal{O}((\mu^{-1}\tau)^{2})+\mathcal{O}((a,b)^{2})\tau}\right\},$
and similarly for the fifth line (recall that $\langle\widetilde{\mathbf{x}}\widetilde{\mathbf{v}}^{\mathsf{T}}\rangle=\mathcal{O}(\mu^{-1})$). The sixth and seventh lines contribute $a[\mathcal{O}(\mu^{-2})+\mathcal{O}((a,b)^{2})]\tau$. Finally, we address the multiplication by $\widetilde{v}^{m}(0)$. The first three lines of Eq. (391) contribute $\mathcal{O}(\mu^{-2}\tau)\mathcal{O}(a,b)$. The fourth and fifth lines contribute $a[\mathcal{O}(\mu^{-2}\tau)+\mathcal{O}(\mu^{-3})+\mathcal{O}((a,b)^{2})\tau]$, while the sixth and seventh lines contribute $a[\mathcal{O}(\mu^{-2})+\mathcal{O}((a,b)^{2})\tau]$.
Now, we analyze the last term in Eq. (391). Similarly as for the drift, we perform a fit for $\tau\lesssim\mathcal{O}(\mu)$:
(393) $\langle[\widetilde{x}^{i},\widetilde{x}^{j}](\tau)\mid\widetilde{\mathbf{x}}(0),\widetilde{\mathbf{v}}(0)\rangle\approx 2\mu^{-1}(\widehat{\mathbf{A}}_{\mathbf{v}}^{-1})^{i}_{i^{\prime}}(\widehat{\mathbf{A}}_{\mathbf{v}}^{-1})^{j}_{j^{\prime}}D^{i^{\prime}j^{\prime}}\tau+2(\breve{b}_{x})^{ij}_{k}\widetilde{x}^{k}(0)+2(\breve{b}_{v})^{ij}_{k}\widetilde{v}^{k}(0).$
Calculating, we obtain:
$(\breve{b}_{x})^{ij}_{k}$ =$=\left\{{(\widehat{\mathbf{A}}_{\mathbf{v}}^{-1})^{i}_{i^{\prime}}(\widehat{\mathbf{A}}_{\mathbf{v}}^{-1})^{j}_{j^{\prime}}\left[{(\widetilde{\mathbf{A}}_{\mathbf{x}})^{-1}\left({\exp({\mu^{-1}\widetilde{\mathbf{A}}_{\mathbf{x}}\tau})-\mathds{1}}\right)}\right]^{k^{\prime}}_{k}(b_{x})^{i^{\prime}j^{\prime}}_{k^{\prime}}+\mathcal{O}(\mu^{-1})b}\right\}$
(394) ×$\quad{}\times[1+\mathcal{O}(\mu)a\mathcal{O}(a,b)],$
(395) $\breve{b}_{v}$ $=\mathcal{O}(\mu^{-1})b[1+\mathcal{O}(\mu)a\mathcal{O}(a,b)]=\mathcal{O}(\mu^{-1})b.$
Note that $\breve{b}_{v}$ computed from Eq. (393) is of the same order as the corresponding quantity computed from the remaining terms in Eq. (391). Now, we again introduce $\tau^{\prime}>0$ and compute:
(396) $\langle([\widetilde{x}^{i},\widetilde{x}^{j}](\tau+\tau^{\prime})-[\widetilde{x}^{i},\widetilde{x}^{j}](\tau^{\prime}))(\widetilde{x}^{k}(\tau^{\prime})-\widetilde{x}^{k}(0))\rangle$
$\qquad{}=(\breve{b}_{x})^{ij}_{l}\langle\widetilde{x}^{l}(\tau^{\prime})(\widetilde{x}^{k}(\tau^{\prime})-\widetilde{x}^{k}(0))\rangle+\mathcal{O}(\mu^{-2})b.$
Again, for $\tau^{\prime}\lesssim\mathcal{O}(\mu)$, the l.h.s. of the above is compared with a magnitude of $\mathcal{O}(\mu^{-2}\tau\tau^{\prime})$. Thus we have shown near-Markovianity with relative error $\mathcal{O}((\tau\tau^{\prime})^{-1})$. This completes the discussion of underdamped systems in the almost-Markovian limit.
@@ 12. Appendix B
Applying Itô’s lemma and taking expectations apparently gives $2\langle x^{2}\rangle=-2\langle x\dot{x}\rangle=1+\langle x^{6}\rangle$. However, the l.h.s. is finite whereas the r.h.s. is infinite. The problem lies in the middle expression. Upon Euler–Maruyama discretization with a fixed time-step, i.e., $x_{i+1}=(1-\Delta t)x_{i}+\sqrt{(1+x_{i}^{6})\Delta t}\cdot\zeta_{i}$ with $\zeta_{i}\sim\mathcal{N}(0,1)$ i.i.d., it would seem that $\langle x_{i}(x_{i+1}-x_{i})/\Delta t\rangle$ should be $-\langle x^{2}\rangle$ because $\langle x^{4}\rangle<\infty$. However, with a fixed time-step $\Delta t$, the sequence $x_{i}$ diverges. To properly simulate this system, an adaptive time-step must be used, i.e., $\Delta t_{i}$ depending on $x_{i}$. We simulated this system using $\Delta t_{i}=0.05(1+x_{i}^{4})^{-1}$ for $10^{8}$ time-steps. Numerically, it appears that $\langle x\dot{x}\rangle=-\infty$ (Fig. 7), in accordance with Eq. (5). Moreover, although the covariance function $\langle x(0)x(\tau)\rangle$ exists for all $\tau$, it is apparently not equal to $\langle x^{2}\rangle e^{-\tau}$ (Fig. 8, calculated using $3\times 10^{5}$ time-steps), as might be expected from the drift function. Thus stochastic force inference fails when using a linear basis function. However, binning reproduces the correct drift (Fig. 9, calculated using $10^{7}$ time-steps). We thus conclude that there is a non-commutation of limits:

### v15

@@ 8.4. Dynamics on long time-scales
(369) $\left\langle\frac{\mathrm{d}}{\mathrm{d}t}\mathbf{f}_{1\widetilde{\mathbf{v}}}(\widetilde{\mathbf{x}},\widetilde{\mathbf{v}})\mid\widetilde{\mathbf{x}},\widetilde{\mathbf{v}}\right\rangle$ $=\mu\mathbf{A}_{\mathbf{v}}[1+\mathcal{O}(\mu^{-2})+a\mathcal{O}(a,b)]\mathbf{f}_{1\widetilde{\mathbf{v}}}(\widetilde{\mathbf{x}},\widetilde{\mathbf{v}}).$
With these expressions in hand, we can proceed similarly to Eq. (171) and obtain, for $\tau>0$:
$\langle\widetilde{\mathbf{x}}(\tau)\mid\widetilde{\mathbf{x}}(0),\widetilde{\mathbf{v}}(0)\rangle$ $=\mathcal{O}(1)\widetilde{\mathbf{x}}(0)+\mathcal{O}(\mu^{-1})a\mathcal{O}(a,b)\widetilde{\mathbf{v}}(0)+\mathcal{O}(a)(\widetilde{x}^{i}(0)\widetilde{x}^{j}(0)-\langle\widetilde{x}^{i}\widetilde{x}^{j}\rangle)$
(370) $\quad{}+\mathcal{O}(\mu^{-1})\mathcal{O}(a)(\widetilde{x}^{i}(0)\widetilde{v}^{j}(0)-\langle\widetilde{x}^{i}\widetilde{v}^{j}\rangle)+\mathcal{O}(\mu^{-1})\mathcal{O}(a)(\widetilde{v}^{i}(0)\widetilde{v}^{j}(0)-\langle\widetilde{v}^{i}\widetilde{v}^{j}\rangle)+\cdots,$
$\langle\widetilde{\mathbf{v}}(\tau)\mid\widetilde{\mathbf{x}}(0),\widetilde{\mathbf{v}}(0)\rangle$ $=\mathcal{O}(1)\widetilde{\mathbf{v}}(0)+a\mathcal{O}(a,b)\widetilde{\mathbf{x}}(0)+\mathcal{O}(\mu^{-1})\mathcal{O}(a)(\widetilde{x}^{i}(0)\widetilde{x}^{j}(0)-\langle\widetilde{x}^{i}\widetilde{x}^{j}\rangle)$
(371) $\quad{}+\mathcal{O}(a)(\widetilde{x}^{i}(0)\widetilde{v}^{j}(0)-\langle\widetilde{x}^{i}\widetilde{v}^{j}\rangle)+\mathcal{O}(a)(\widetilde{v}^{i}\widetilde{v}^{j}-\langle\widetilde{v}^{i}(0)\widetilde{v}^{j}(0)\rangle)+\cdots.$
We now introduce another time-lag $\tau^{\prime}>0$ and consider the quantities:
(372) $\langle x^{i}(\tau+\tau^{\prime})x^{j}(\tau^{\prime})\rangle,\quad\langle x^{i}(\tau+\tau^{\prime})(x^{j}(\tau^{\prime})-x^{j}(0))\rangle.$
For $\tau,\tau^{\prime}\sim\mathcal{O}(\mu)$, the above quantities vary by $\mathcal{O}(1)$. We translate Eqs. (370)–(371) in time by $\tau^{\prime}$ and apply the law of iterated expectations. We see that all terms in $\mathbf{x}(\tau^{\prime})$ except for $\widetilde{\mathbf{x}}(\tau^{\prime})$ contribute only $\mathcal{O}(\mu^{-2})$. We now consider the quantities:
(373) $\langle x^{i}(\tau+\tau^{\prime})x^{j}(\tau^{\prime})x^{k}(\tau^{\prime})\rangle,\quad\langle x^{i}(\tau+\tau^{\prime})x^{j}(\tau^{\prime})(x^{k}(\tau^{\prime})-x^{k}(0))\rangle,\quad\langle x^{i}(\tau+\tau^{\prime})(x^{j}(\tau^{\prime})-x^{j}(0))(x^{k}(\tau^{\prime})-x^{k}(0))\rangle.$
We seek to show that all terms in $\mathbf{x}(\tau^{\prime})$ involving $\widetilde{\mathbf{v}}(\tau^{\prime})$ contribute $\mathcal{O}(\mu^{-2})\mathcal{O}(a,b)$ to the above quantities. For this, we need to use:
(374) $\langle\widetilde{v}^{i}\widetilde{x}^{j}\widetilde{x}^{k}\widetilde{x}^{l}\rangle$ $=\mathcal{O}(\mu^{-1}),$
(375) $\langle\widetilde{v}^{i}(\tau^{\prime})\widetilde{x}^{j}(0)\widetilde{x}^{k}(0)\widetilde{x}^{l}(0)\rangle$ $=\mathcal{O}(\mu^{-1}),$
and similarly whenever any subset of the factors $\widetilde{\mathbf{x}}(0)$ is replaced by $\widetilde{\mathbf{x}}(\tau^{\prime})$. Also,
(376) $\langle(\widetilde{v}^{i}\widetilde{v}^{j}-\langle\widetilde{v}^{i}\widetilde{v}^{j}\rangle)\widetilde{x}^{k}\widetilde{x}^{l}\rangle$ $=\mathcal{O}(\mu^{-2})+\mathcal{O}((a,b)^{2}),$
(377) $\langle(\widetilde{v}^{i}(\tau^{\prime})\widetilde{v}^{j}(\tau^{\prime})-\langle\widetilde{v}^{i}\widetilde{v}^{j}\rangle)\widetilde{x}^{k}(0)\widetilde{x}^{l}(0)\rangle$ $=O(\mu^{-2})+\mathcal{O}((a,b)^{2}),$
and similarly when either factor $\widetilde{\mathbf{x}}(0)$ is replaced by $\widetilde{\mathbf{x}}(\tau^{\prime})$. We have thus established that on $\mathcal{O}(\mu)$ time-scales, the conditional expectation of $\mathbf{x}$ is Markovian with $\mathcal{O}(\mu^{-2})$ error. Higher-order Hermite polynomials (e.g. Eq. (172)) can be handled similarly.
@@ 12. Appendix B
Applying Itô’s lemma and taking expectations apparently gives $2\langle x^{2}\rangle=-2\langle x\dot{x}\rangle=1+\langle x^{6}\rangle$. However, the l.h.s. is finite whereas the r.h.s. is infinite [^24] The same problem is encountered if $x^{6}$ is replaced by $x^{4}$ in Eq. (399).. The problem lies in the middle expression. Upon Euler–Maruyama discretization with a fixed time-step, i.e., $x_{i+1}=(1-\Delta t)x_{i}+\sqrt{(1+x_{i}^{6})\Delta t}\cdot\zeta_{i}$ with $\zeta_{i}\sim\mathcal{N}(0,1)$ i.i.d., it would seem that $\langle x_{i}(x_{i+1}-x_{i})/\Delta t\rangle$ should be $-\langle x^{2}\rangle$ because $\langle x^{4}\rangle<\infty$. However, with a fixed time-step $\Delta t$, the sequence $x_{i}$ diverges. To properly simulate this system, an adaptive time-step must be used, i.e., $\Delta t_{i}$ depending on $x_{i}$. We simulated this system using $\Delta t_{i}=0.05(1+x_{i}^{4})^{-1}$ for $10^{8}$ time-steps. Numerically, it appears that $\langle x\dot{x}\rangle=-\infty$ (Fig. 7), in accordance with Eq. (5). Moreover, although the covariance function $\langle x(0)x(\tau)\rangle$ exists for all $\tau$, it is apparently not equal to $\langle x^{2}\rangle e^{-\tau}$ (Fig. 8, calculated using $3\times 10^{5}$ time-steps), as might be expected from the drift function. Thus stochastic force inference fails when using a linear basis function. However, binning reproduces the correct drift (Fig. 9, calculated using $10^{7}$ time-steps). We thus conclude that there is a non-commutation of limits: